OpenStudy (anonymous):

how do i find out the mixed number? Which of these is a mixed number? 7 3 2/7ths 11/4ths 1/9th

jimthompson5910 (jim_thompson5910):

Mixed numbers look like this \[\Large 3 \frac{1}{4}\] or \[\Large 9 \frac{7}{12}\]

OpenStudy (anonymous):

so its 3 2/7ths

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

\[\Large 3 \frac{2}{7}\]

OpenStudy (anonymous):

oh ok so a mixed number has a full number and the fraction

jimthompson5910 (jim_thompson5910):

yes, the basic form is \[\Large \text{Whole}\frac{\text{numerator}}{\text{denominator}}\]

OpenStudy (anonymous):

oh ok well thanks for the help sorry it took me a while to reply

jimthompson5910 (jim_thompson5910):

it's ok, yw